OpenStudy (anonymous):

The perimeter of a rectangle is 96 ft. The ratio of its length to its width is 7 : 5. What are the dimensions of the rectangle? @chaser71 A.30 ft by 18 ft B.28 ft by 20 ft C.24 ft by 24 ft D.22 ft by 26 ft

OpenStudy (anonymous):

B

OpenStudy (anonymous):

2a + 2b = 96 a / b = 7 / 5 2a + 2b = 96 a + b = 48 a = 48 - b (48 - b) / b = 7 / 5 7b = 5(48 - b) 7b = 240 - 5b 12b = 240 b = 20 2a + 2(20) = 96 2a + 40 = 96 2a = 56 a = 28

OpenStudy (anonymous):

can you write up a testimonial for me? :)

OpenStudy (anonymous):

thanks :) andd sure :p hhah